Job-Order Cost Sheets, Balance in Work in Process and Finished Goods

Prull Company, a job-order costing firm, worked on three jobs in July. Data are as follows:

Job 86

Job 87

Job 88

Balance, July 1 $21,320 $6,300 $0
Direct materials $10,490 $12,280 $16,210
Direct labor $16,000 $12,150 $23,990
Machine hours 480 290 1,020

Overhead is applied to jobs at the rate of $16 per machine hour. By July 31, Jobs 86 and 88 were completed. Jobs 82 and 86 were sold. Job 87 remained in process. On July 1, the balance in Finished Goods was $48,900 (consisting of Job 82 for $25,550 and Job 84 for $23,350).

Prull prices its jobs at cost plus 30%. During July, variable marketing expenses were 3% of sales, and fixed marketing expenses were $1,990; administrative expenses were $4,780. (Round all amounts to the nearest dollar.)

Required:
1. Prepare job-order cost sheets for all jobs in process during July, showing all costs through July 31.
2. Calculate the balance in Work in Process on July 31.
3. Calculate the balance in Finished Goods on July 31.
4. Calculate Cost of Goods Sold for July.
5. Calculate operating income for Prull Company for the month of July.

Job-Order Cost Sheets

1. Prepare job-order cost sheets for all jobs in process during July, showing all costs through July 31. All amount cells must have an entry; enter zeros as needed.

Prull Company

Job-Order Cost Sheets

For the Month of July

1

Job 86

Job 87

Job 88

2

Balance, July 1

3

Direct materials

4

Direct labor

5

Applied overhead

6

Total

Balances and Calculations

2. Calculate the balance in Work in Process on July 31.

The balance in Work in Process on July 31 is .

3. Calculate the balance in Finished Goods on July 31.

The balance in Finished Goods on July 31 is .

4. Calculate Cost of Goods Sold for July.

The cost of goods sold for July is .

5. Calculate operating income for Prull Company for the month of July.

The operating income for the month of July is .
Answer #1

1. Prepare job-order cost sheets for all jobs in process during July, showing all costs through July 31. All amount cells must have an entry; enter zeros as needed.

Prull Company

Job-Order Cost Sheets

For the Month of July

1

Job 86

Job 87

Job 88

2

Balance, July 1

21320 6300 0

3

Direct materials

10490 12280 16210

4

Direct labor

16000 12150 23990

5

Applied overhead

480*16 = 7680 290*16 = 4640 1020*16 = 16320

6

Total

55490 35370 56520

Balances and Calculations

2. Calculate the balance in Work in Process on July 31.

The balance in Work in Process on July 31 is 35370

3. Calculate the balance in Finished Goods on July 31.

The balance in Finished Goods on July 31 is 56520

4. Calculate Cost of Goods Sold for July.

The cost of goods sold for July is 25550+55490 = 81040

5. Calculate operating income for Prull Company for the month of July.

Sales (81040*1.3) 105352
Cost of goods sold -81040
Variable marketing (81040*3%) -2431
Fixed marketing expense -1990
Administrative expense -4780
Operating income 15111